OpenStudy (anonymous):

0.7(3x+4).>=1.3-(x+2)

OpenStudy (anonymous):

2,1x+1,28 >= 1,3-x-2 2,1x+x>=-1,28-0,7 3,2x>=0,58 x>=0,18

OpenStudy (anonymous):

Solution: x\[>=\]-35\/31